OpenStudy (anonymous):

When you simplify each side of the equation 2(3n – 5) + 4 = 15 – (9n – 4), you will get the equivalent equation 6n – 6 = -9n + 11. True False

OpenStudy (anonymous):

2(3n – 5) + 4 = 15 – (9n – 4) 6n-10+4=15-9n+4